# Mobile-Efficient Speech Emotion Recognition Using DistilHuBERT: A Cross-Corpus Validation Study

### v1

@@ Abstract
Speech Emotion Recognition (SER) has significant potential for mobile applications, yet deployment remains constrained by the computational demands of state-of-the-art transformer architectures. This paper presents a mobile-efficient SER system based on DistilHuBERT, a distilled and 8-bit quantized transformer that achieves 92% parameter reduction compared to full-scale Wav2Vec 2.0 models while maintaining competitive accuracy. We conduct a rigorous 5-fold Leave-One-Session-Out (LOSO) cross-validation on the IEMOCAP dataset to ensure speaker independence, augmented with cross-corpus training on CREMA-D to enhance generalization. Cross-corpus training with CREMA-D yields a 1.2% improvement in Weighted Accuracy, a 1.4% gain in Macro F1-score, and a 32% reduction in cross-fold variance, with the Neutral class showing the most substantial benefit at 5.4% F1-score improvement. Our approach achieves an Unweighted Accuracy of 61.4% with a quantized model footprint of only 23 MB, representing approximately 91% of full-scale baseline performance. Cross-corpus evaluation on RAVDESS reveals that the theatrical nature of acted emotions causes predictions to cluster by arousal level rather than valence—happiness is systematically confused with anger due to acoustic saturation in high-energy expressions. Despite this theatricality effect reducing overall RAVDESS accuracy to 43.29%, the model maintains robust arousal detection with 97% recall for anger and 64% for sadness. These findings establish a Pareto-optimal tradeoff between model size and accuracy, enabling practical affect recognition on resource-constrained mobile devices.
@@ II Introduction
Speech Emotion Recognition (SER) is critical for responsive mobile applications, including mental health monitoring, customer service analytics, and adaptive human-computer interaction. The ability to automatically detect emotional states from vocal cues enables systems to provide personalized feedback, detect distress signals, and enhance user experience through affectively aware interfaces. However, the deployment of SER systems on mobile devices faces a fundamental challenge: state-of-the-art models based on self-supervised transformer architectures such as Wav2Vec 2.0 [8] and HuBERT [5] achieve high accuracy but require hundreds of megabytes of storage and substantial computational resources, making them impractical for edge deployment.
@@
Beyond architectural efficiency, a second critical challenge in SER research is generalization. Many published results rely on cross-validation protocols that inadvertently introduce speaker leakage, where the same speaker appears in both training and test sets. This methodological flaw leads to inflated accuracy estimates, as models learn speaker-specific vocal traits rather than generalizable emotional patterns. To address this, we adopt a strict 5-fold Leave-One-Session-Out (LOSO) protocol on the IEMOCAP dataset [1], ensuring that each test fold contains entirely unseen speakers. This rigorous evaluation standard, while resulting in lower absolute accuracy, provides a more realistic estimate of deployment performance.
@@
This paper makes the following contributions: (1) We demonstrate that DistilHuBERT achieves 91% of full-scale Wav2Vec 2.0 performance while reducing model size by 92%, establishing a new Pareto frontier for mobile SER. (2) We conduct a methodologically rigorous 5-fold LOSO evaluation, avoiding the speaker leakage issues prevalent in prior work. (3) We quantify the regularization benefits of cross-corpus training with CREMA-D, showing a 1.2% gain in Weighted Accuracy and a 32% reduction in cross-fold variance. (4) We provide detailed error analysis on RAVDESS [7], revealing that the model maintains robust arousal detection (97% recall for anger) despite domain shift, with errors systematically clustered by energy level rather than random misclassification. (5) We propose a practical deployment architecture utilizing Voice Activity Detection and temporal aggregation to mitigate neutral-class bias in long-form voice note analysis.
@@ III Methodology
In this section, we detail our approach to Speech Emotion Recognition (SER), emphasizing a pipeline designed for mobile deployment and cross-corpus robustness. Our methodology is characterized by a multi-stage acoustic workflow, the use of a distilled transformer architecture, and a specialized training objective to handle class imbalance.
@@ III-A Experimental Framework and Dataset Integration
We utilize the IEMOCAP dataset [1] as our primary benchmark, focusing on four categorical emotions: anger, happiness (including excitement), neutral, and sadness. To evaluate speaker independence, we implement a strict 5-fold Leave-One-Session-Out (LOSO) cross-validation, following the protocol of Pepino et al. [2]. In each fold, the model is trained on three sessions, validated on one, and tested on the remaining session.
To enhance robustness, we incorporate the CREMA-D dataset [6] as a permanent training component. In adherence to rigorous evaluation standards, CREMA-D samples are exclusively used for training and are never included in validation or test sets. This configuration ensures that our performance metrics reflect true generalization to unseen IEMOCAP speakers across distinct recording environments.
@@ III-C Feature Extraction and Architecture Selection
For feature extraction, we prioritize mobile-deployability through the use of DistilHuBERT [4]. While standard Wav2Vec 2.0 models [2] offer high performance, their parameter count is often prohibitive for edge devices. DistilHuBERT utilizes knowledge distillation to compress the Hidden-unit BERT (HuBERT) architecture into a shallower transformer encoder. Our choice is motivated by hardware efficiency; upon 8-bit quantization, the model footprint is approximately 23 MB. In our implementation, the feature encoder remains frozen to leverage pre-trained acoustic representations, while the downstream classification head is fine-tuned for categorical emotion recognition.
@@ III-D Optimization Strategy and Hyperparameters
To address the class imbalance detailed in Table II, we utilize an Adaptive Focal Loss function. This objective integrates balanced class weights ($\alpha_{t}$) and a focusing parameter ($\gamma=2.0$) to prioritize difficult, under-represented samples. The loss function is further regularized with a label smoothing factor of 0.1, defined as:
@@
IV Results and Discussion
TABLE I: Performance Comparison of 5-Fold LOSO Cross-Validation on IEMOCAP
@@ IV Results and Discussion
TABLE II: Distribution of Emotional Classes Across IEMOCAP and CREMA-D
Dataset Anger Hap/Exc Neutral Sad Total
IEMOCAP 1,103 1,636 1,708 1,084 5,531
CREMA-D (Train) 1,271 1,271 1,087 1,271 4,900
Combined Total 2,374 2,907 2,795 2,355 10,431
Avg. Class Weight 1.065 0.913 0.974 1.070
The experimental results for the 5-fold LOSO cross-validation on the IEMOCAP dataset are detailed in Table I. Our analysis examines the trade-offs between architectural efficiency, the regularization effects of multi-corpus training, and the methodological rigor required for real-world deployment.
@@ IV-A Cross-Corpus Regularization via CREMA-D
A primary focus of this study was the integration of the CREMA-D dataset as a permanent training anchor. As seen in Table II, the inclusion of 4,900 additional samples nearly doubled the available training data. This strategy yielded a 1.2% increase in Weighted Accuracy (WA) and a 1.4% improvement in Macro F1-score. More critically, the standard deviation of the WA across the five folds decreased from $0.031$ to $0.021$, representing a 32% reduction in cross-fold variance. This substantial decrease in performance variability is a strong indicator that CREMA-D functions as an effective regularization mechanism. The acoustic diversity provided by the 91 additional speakers in CREMA-D—each with distinct vocal tract characteristics, recording equipment, and expressive styles—serves as a stabilizing regularizer, reducing the model’s sensitivity to the idiosyncratic vocal characteristics of individual IEMOCAP sessions. Rather than merely increasing dataset size, the introduction of cross-corpus heterogeneity prevents the model from overfitting to corpus-specific artifacts such as room acoustics, microphone frequency response, or the limited phonetic coverage of a single speaker pool.
@@
The class-wise analysis reveals a 5.4% gain in the Neutral class F1-score, suggesting that multi-corpus exposure helps the model calibrate a more robust baseline for vocal energy. However, we observed a 2.9% decrease in Sadness recall. This discrepancy points to a ”theatricality gap”: CREMA-D’s sadness is characterized by high-intensity, acted prosody, whereas IEMOCAP features naturalistic, low-intensity sadness. The model’s exposure to the theatrical variant appears to have shifted the decision boundary, leading to increased confusion with neutral speech in the test set. Despite this, the overall gain in F1-score and the dramatic improvement in cross-fold stability justify the use of CREMA-D for broader acoustic coverage and robust generalization.
IV-B Comparative Benchmarking and Methodological Rigor
@@
To evaluate the proposed DistilHuBERT approach, we contextualize our findings against both high-capacity and efficiency-oriented benchmarks. First, we consider the work of Pepino et al. [2], who utilized a full Wav2Vec 2.0 Base model (approx. 318 MB) to achieve an Unweighted Accuracy (UA) of 67.2%. Our distilled implementation achieved a UA of 61.4%, representing approximately 91% of the performance of the full-scale baseline while utilizing a 23 MB quantized footprint. This 92% reduction in parameter size is critical for mobile deployment, where memory and thermal constraints are paramount.
@@
Furthermore, we address the performance of the LIGHT-SERNET architecture by Aftab et al. [3], which reported a UA of 70.76%. While their absolute accuracy is higher, there is a significant methodological distinction: their study utilized a 10-fold cross-validation protocol. In the context of SER, standard k-fold CV often involves a random shuffle of the dataset, which is susceptible to ”speaker leakage.” In such cases, the model may inadvertently learn speaker-specific biological identities rather than generalized emotional cues. By contrast, our adherence to a strict 5-fold LOSO protocol ensures the model generalizes to entirely unseen speakers. This rigor accounts for the lower absolute UA but confirms the model’s viability for deployment with unknown users.
@@ IV-C Architectural Capacity and the Pareto Frontier
The results suggest that our implementation may be approaching the theoretical capacity limit for a two-layer distilled transformer on the SER task. Emotion recognition requires the simultaneous extraction of low-level prosody and high-level temporal dependencies. The reduced depth of DistilHuBERT, while efficient, likely limits the model’s ability to resolve the most subtle emotional nuances found in naturalistic interactions.
However, the efficacy of the Adaptive Focal Loss (weights in Table II) ensured the model did not collapse into majority-class prediction strategies. Our system thus represents a ”Pareto-optimal” point—maximizing accuracy per megabyte of model weight. This confirms that a carefully regularized, distilled model can provide reliable affect recognition within the strict envelopes of edge hardware, bridging the gap between laboratory performance and real-world mobile feasibility.
V Production Model Development and Cross-Corpus Robustness
@@ V-B Generalization and the Theatricality Effect in Error Analysis
The robustness of the production model was scrutinized through a cross-corpus evaluation on the RAVDESS dataset. While the domain shift between the semi-spontaneous IEMOCAP data and the theatrical RAVDESS expressions resulted in an overall accuracy of 43.29%, the model demonstrated a significant capability to identify underlying physiological arousal states.
As illustrated in Fig. 1, the model’s errors are systematically clustered by arousal levels rather than random distribution. We hypothesize that this phenomenon is a direct result of the theatrical nature of the RAVDESS corpus. In RAVDESS, actors prioritize clarity and communicative intent, leading to ”over-acted” expressions that emphasize vocal intensity and pitch range at the expense of subtlety. For a model trained on the more nuanced, semi-spontaneous features of IEMOCAP, this theatricality creates an acoustic ”bleeding” effect. Specifically, high-energy samples (Happiness and Anger) become acoustically saturated, causing the model to lose the subtle spectral distinctions required to differentiate valence; consequently, it defaults to the more acoustically dominant ”Anger” class. Similarly, the low-energy theatricality of ”Neutral” samples often mimics the prosodic markers of ”Sadness,” leading to a significant recall drop for the neutral class. Despite this valence confusion, the model remains a highly robust detector of arousal intensity, maintaining a 97% recall for Anger and a 64% recall for Sadness across disparate recording environments.
Figure 1: Cross-corpus confusion matrix on RAVDESS. The results highlight the model’s tendency to cluster high-arousal (Anger/Happy) and low-arousal (Neutral/Sad) states, suggesting that theatrical over-emphasis in the target corpus obscures the subtle valence cues learned during training. Overall Accuracy: 43.29%; Unweighted Recall: 46.18%.
@@ V-C Proposed Deployment Architecture
Given the model’s high sensitivity to high-arousal and negative-valence states, we propose a specialized deployment pipeline for processing long-form voice notes. This architecture initiates with a Voice Activity Detection (VAD) layer to eliminate non-speech segments, thereby mitigating ”Neutral” bias during inference. The filtered audio is then segmented into 8-second windows to capture necessary macro-prosodic contours. Finally, the system employs an aggregated averaging mechanism across all windows within a voice note. This pooling strategy serves as a temporal filter, smoothing out momentary classification jitter and providing a stable emotional sentiment for the entire communication. By focusing on longitudinal emotional trends, this pipeline leverages the model’s strengths in identifying anger and sadness while compensating for the cross-corpus precision drop in positive affect caused by domain-specific theatricality.

### v2

# Distilled HuBERT for Mobile Speech Emotion Recognition: A Cross-Corpus Validation Study

@@ Abstract
Speech Emotion Recognition (SER) has significant potential for mobile applications, yet deployment remains constrained by the computational demands of state-of-the-art transformer architectures. This paper presents a mobile-efficient SER system based on DistilHuBERT, a distilled and 8-bit quantized transformer that achieves approximately 92% parameter reduction compared to full-scale Wav2Vec 2.0 models while maintaining competitive accuracy. We conduct a rigorous 5-fold Leave-One-Session-Out (LOSO) cross-validation on the IEMOCAP dataset to ensure speaker independence, augmented with cross-corpus training on CREMA-D to enhance generalization. Cross-corpus training with CREMA-D yields a 1.2% improvement in Weighted Accuracy, a 1.4% gain in Macro F1-score, and a 32% reduction in cross-fold variance, with the Neutral class showing the most substantial benefit at 5.4% F1-score improvement. Our approach achieves an Unweighted Accuracy of 61.4% with a quantized model footprint of only 23 MB, representing approximately 91% of the Unweighted Accuracy of a full-scale baseline. Cross-corpus evaluation on RAVDESS reveals that the theatrical nature of acted emotions causes predictions to cluster by arousal level rather than by specific emotion categories—happiness predictions systematically bleed into anger predictions, and sadness predictions bleed into neutral predictions, due to acoustic saturation when actors prioritize clarity over subtlety. Despite this theatricality effect reducing overall RAVDESS accuracy to 46.64%, the model maintains robust arousal detection with 99% recall for anger, 55% recall for neutral, and 27% recall for sadness. These findings demonstrate a Pareto-optimal tradeoff between model size and accuracy, enabling practical affect recognition on resource-constrained mobile devices.
@@ II Introduction
Speech Emotion Recognition (SER) is critical for responsive mobile applications, including mental health monitoring, customer service analytics, and adaptive human-computer interaction. The ability to automatically detect emotional states from vocal cues enables systems to provide personalized feedback, detect distress signals, and enhance user experience through affectively aware interfaces. However, the deployment of SER systems on mobile devices faces a fundamental challenge: state-of-the-art models based on self-supervised transformer architectures such as Wav2Vec 2.0 [8] and Hidden-unit BERT (HuBERT) [5] achieve high accuracy but require hundreds of megabytes of storage and substantial computational resources, making them impractical for edge deployment.
@@
Beyond architectural efficiency, a second critical challenge in SER research is generalization. Many published results rely on cross-validation protocols that may introduce speaker leakage, where the same speaker appears in both training and test sets. This methodological limitation can lead to inflated accuracy estimates, as models may learn speaker-specific vocal traits rather than generalizable emotional patterns. To address this, we adopt a strict 5-fold Leave-One-Session-Out (LOSO) protocol on the IEMOCAP dataset [1], ensuring that each test fold contains entirely unseen speakers. This rigorous evaluation standard, while resulting in lower absolute accuracy, provides a more realistic estimate of deployment performance.
@@
This paper makes the following contributions: (1) We demonstrate that DistilHuBERT achieves 91% of full-scale Wav2Vec 2.0 performance while reducing model size by 92%, approaching a Pareto frontier for mobile SER. (2) We conduct a methodologically rigorous 5-fold LOSO evaluation, avoiding the speaker leakage issues prevalent in prior work. (3) We quantify the regularization benefits of cross-corpus training with CREMA-D, showing a 1.2% gain in Weighted Accuracy and a 32% reduction in cross-fold variance. (4) We provide detailed error analysis on RAVDESS [7], revealing that the model maintains robust arousal detection (99% recall for anger, 55% for neutral, 27% for sadness) despite domain shift, with errors systematically clustered by energy level rather than random misclassification. (5) We propose a practical deployment architecture utilizing Voice Activity Detection and temporal aggregation to mitigate neutral-class bias in long-form voice note analysis.
@@ III Methodology
In this section, we detail our approach to SER, emphasizing a pipeline designed for mobile deployment and cross-corpus robustness. Our methodology is characterized by a multi-stage acoustic workflow, the use of a distilled transformer architecture, and a specialized training objective to handle class imbalance.
@@ III-A Experimental Framework and Dataset Integration
We utilize the IEMOCAP dataset [1] as our primary benchmark, focusing on four categorical emotions: anger, happiness (including excitement), neutral, and sadness. To evaluate speaker independence, we implement a strict 5-fold LOSO cross-validation, following the protocol of Pepino et al. [2]. In each fold, the model is trained on three sessions, validated on one, and tested on the remaining session.
To enhance robustness, we incorporate the CREMA-D dataset [6] as a permanent training component. In adherence to rigorous evaluation standards, CREMA-D samples are exclusively used for training and are never included in validation or test sets. This configuration helps ensure that our performance metrics reflect generalization to unseen IEMOCAP speakers across distinct recording environments.
@@ III-C Feature Extraction and Architecture Selection
For feature extraction, we prioritize mobile-deployability through the use of DistilHuBERT [4]. While standard Wav2Vec 2.0 models [2] offer high performance, their parameter count is often prohibitive for edge devices. DistilHuBERT utilizes knowledge distillation to compress the HuBERT architecture into a shallower transformer encoder. Our choice is motivated by hardware efficiency; upon 8-bit quantization, the model footprint is approximately 23 MB. In our implementation, the feature encoder remains frozen to leverage pre-trained acoustic representations, while the downstream classification head is fine-tuned for categorical emotion recognition.
@@ III-D Optimization Strategy and Hyperparameters
To address the class imbalance detailed in Table I, we utilize an Adaptive Focal Loss function. This objective integrates balanced class weights ($\alpha_{t}$) and a focusing parameter ($\gamma=2.0$) to prioritize difficult, under-represented samples. The loss function is further regularized with a label smoothing factor of 0.1, defined as:
@@
IV Results
TABLE I: Distribution of Emotional Classes Across IEMOCAP and CREMA-D Datasets
Dataset Anger Hap/Exc Neutral Sad Total
IEMOCAP 1,103 1,636 1,708 1,084 5,531
CREMA-D (Train) 1,271 1,271 1,087 1,271 4,900
Combined Total 2,374 2,907 2,795 2,355 10,431
Avg. Class Weight 1.065 0.913 0.974 1.070
TABLE II: Performance Comparison of 5-Fold LOSO Cross-Validation on IEMOCAP Dataset
@@ IV Results
The experimental results for the 5-fold LOSO cross-validation on the IEMOCAP dataset are detailed in Table II. Our analysis examines the trade-offs between architectural efficiency, the regularization effects of multi-corpus training, and the methodological rigor required for real-world deployment.
@@ IV-A Cross-Corpus Regularization via CREMA-D
A primary focus of this study was the integration of the CREMA-D dataset as a permanent training anchor. As seen in Table I, the inclusion of 4,900 additional samples nearly doubled the available training data. This strategy yielded modest improvements: a 1.2% increase in Weighted Accuracy (WA) and a 1.4% improvement in Macro F1-score. More critically, the standard deviation of the WA across the five folds decreased from $0.031$ to $0.021$, representing a 32% reduction in cross-fold variance. The relatively modest accuracy gains despite doubling the training data suggest that the model may be approaching its capacity limits for this task, as discussed in Section III-C. However, the substantial decrease in performance variability indicates that CREMA-D functions primarily as an effective regularization mechanism rather than simply providing more training examples. The acoustic diversity provided by the 91 additional speakers in CREMA-D—each with distinct vocal tract characteristics, recording equipment, and expressive styles—serves as a stabilizing regularizer, reducing the model’s sensitivity to the idiosyncratic vocal characteristics of individual IEMOCAP sessions. Rather than merely increasing dataset size, the introduction of cross-corpus heterogeneity helps prevent the model from overfitting to corpus-specific artifacts such as room acoustics, microphone frequency response, or the limited phonetic coverage of a single speaker pool.
@@
The class-wise analysis reveals a 5.4% gain in the Neutral class F1-score, suggesting that multi-corpus exposure helps the model calibrate a more robust baseline for vocal energy. However, we observed a 2.9% decrease in Sadness recall. This discrepancy points to a ”theatricality gap”: CREMA-D’s sadness is characterized by high-intensity, acted prosody, whereas IEMOCAP features naturalistic, low-intensity sadness. The model’s exposure to the theatrical variant appears to have shifted the decision boundary, leading to increased confusion with neutral speech in the test set. Despite this, the overall gain in F1-score and the substantial improvement in cross-fold stability justify the use of CREMA-D for broader acoustic coverage and robust generalization.
IV-B Comparative Benchmarking
@@
To evaluate the proposed DistilHuBERT approach, we contextualize our findings against both high-capacity and efficiency-oriented benchmarks. First, we consider the work of Pepino et al. [2], who utilized a full Wav2Vec 2.0 Base model (approx. 318 MB) to achieve an UA of 67.2%. Our distilled implementation achieved a UA of 61.4%, representing approximately 91% of the performance of the full-scale baseline while utilizing a 23 MB quantized footprint. This 92% reduction in parameter size is critical for mobile deployment, where memory and thermal constraints are paramount.
@@
Furthermore, we address the performance of the LIGHT-SERNET architecture by Aftab et al. [3], which reported a UA of 70.76%. While their absolute accuracy is higher, there is a significant methodological distinction: their study utilized a 10-fold cross-validation protocol. In the context of SER, standard k-fold CV often involves a random shuffle of the dataset, which is susceptible to ”speaker leakage.” In such cases, the model may inadvertently learn speaker-specific biological identities rather than generalized emotional cues. By contrast, our adherence to a strict 5-fold LOSO protocol helps ensure the model generalizes to entirely unseen speakers. This rigor accounts for the lower absolute UA but demonstrates the model’s viability for deployment with unknown users.
@@ IV-C Architectural Capacity and the Pareto Frontier
The results suggest that our implementation may be approaching the theoretical capacity limit for a two-layer distilled transformer on the SER task. This hypothesis is supported by the observation that doubling the training data (via CREMA-D integration) yielded only modest accuracy improvements (1.2% WA, 1.4% F1) despite substantially reducing cross-fold variance. Such diminishing returns are characteristic of models operating near their representational capacity. Emotion recognition requires the simultaneous extraction of low-level prosody and high-level temporal dependencies. The reduced depth of DistilHuBERT, while efficient, likely limits the model’s ability to resolve the most subtle emotional nuances found in naturalistic interactions.
However, the efficacy of the Adaptive Focal Loss (weights in Table I) ensured the model did not collapse into majority-class prediction strategies. Our system thus approaches a ”Pareto-optimal” point—maximizing accuracy per megabyte of model weight. This demonstrates that a carefully regularized, distilled model can provide reliable affect recognition within the strict envelopes of edge hardware, bridging the gap between laboratory performance and real-world mobile feasibility.
V Discussion
@@ V-B Generalization and the Theatricality Effect in Error Analysis
The robustness of the production model was scrutinized through a cross-corpus evaluation on the RAVDESS dataset. While the domain shift between the semi-spontaneous IEMOCAP data and the theatrical RAVDESS expressions resulted in an overall accuracy of 46.64%, the model demonstrated a significant capability to identify underlying physiological arousal states.
As illustrated in Fig. 1, the model’s predictions cluster by arousal level rather than by specific emotion categories. We hypothesize that this phenomenon is likely a result of the theatrical nature of the RAVDESS corpus. In RAVDESS, actors prioritize clarity and communicative intent, leading to ”over-acted” expressions that emphasize vocal intensity and pitch range at the expense of subtlety. For a model trained on the more nuanced, semi-spontaneous features of IEMOCAP, this theatricality creates an acoustic ”bleeding” effect where predictions collapse into arousal-based clusters. Specifically, happiness predictions systematically bleed into anger predictions, as both high-arousal states become acoustically saturated, causing the model to lose the subtle spectral distinctions required to differentiate valence; consequently, it defaults to the more acoustically dominant ”Anger” class. Similarly, sadness predictions bleed into neutral predictions, as the low-energy theatricality of both states leads the model to cluster them together rather than distinguishing their valence differences. This arousal-based clustering indicates that the model maintains robust detection of physiological arousal states despite the domain shift, maintaining a 99% recall for Anger, a 55% recall for Neutral, and a 27% recall for Sadness across disparate recording environments.
Figure 1: Cross-corpus confusion matrix on RAVDESS. The results highlight the model’s tendency to cluster high-arousal (Anger/Happy) and low-arousal (Neutral/Sad) states, suggesting that theatrical over-emphasis in the target corpus obscures the subtle valence cues learned during training. Overall Accuracy: 46.64%; Unweighted Recall: 45.57%.
@@ V-C Proposed Deployment Architecture
Given the model’s high sensitivity to high-arousal states (particularly anger), we propose a specialized deployment pipeline for processing long-form voice notes. This architecture initiates with a Voice Activity Detection layer to eliminate non-speech segments, thereby mitigating ”Neutral” bias during inference. The filtered audio is then segmented into 8-second windows to capture necessary macro-prosodic contours. Finally, the system employs an aggregated averaging mechanism across all windows within a voice note. This pooling strategy serves as a temporal filter, smoothing out momentary classification jitter and providing a stable emotional sentiment for the entire communication. By focusing on longitudinal emotional trends, this pipeline leverages the model’s strength in identifying high-arousal states (particularly anger) while compensating for the cross-corpus precision drop in positive affect caused by domain-specific theatricality.